### v1

# Past, Present, and Future of Citation Practices in HCI

Abstract.
Science is a complex system comprised of many scientists who individually make collective decisions that, due to the size and nature of the academic system, largely do not affect the system as a whole. However, certain decisions at the meso-level of research communities, such as the Human-Computer Interaction (HCI) community, may result in deep and long-lasting behavioral changes in scientists. In this article, we provide evidence on how a change in editorial policies introduced at the ACM CHI Conference in 2016 launched the CHI community on an expansive path, denoted by a year-by-year increase in the mean number of references included in CHI articles. If this near-linear trend continues undisrupted, an article in CHI 2030 will include on average almost 130 references. Our meta-research provides insights into how the nature and meaning of citation practices in HCI have changed, influenced by factors such as digital accessibility of resources and academic pressures. The observed trend towards more citations reflects a citation culture where quantity is prioritized over quality, contributing to both author and peer reviewer fatigue. This article underscores the value of meta-research for research communities and the profound impact that meso-level policy adjustments have on the evolution of scientific fields and disciplines, urging stakeholders to carefully consider the broader implications of such changes.
journalyear: 2024
Figure 1. Number of references per article in the Proceedings of the ACM CHI Conference from 1981 to 2024.
[…]
Number of references per article in the Proceedings of the ACM CHI Conference from 1981 to 2024.
1. Introduction
Citation serves as the backbone of academic rigor, enabling researchers to build upon previous work, acknowledge contributions, and weave a rich tapestry of interdisciplinary knowledge. Editorial policies shape the landscape of academic publishing, influencing not only the structure and content of the research published but also the citation practices of researchers. Citation practices, as a tangible expression of scholarly discourse, provide valuable insights into the evolving norms and values of the CHI community. Understanding the impact of policy decisions on citation practices is crucial for ensuring that scholarly communication serves both the authors and the broader field effectively.
In this work, we examine how the citation practices of researchers in the Human-Computer Interaction (HCI) community have changed after a significant editorial policy decision introduced in 2016. In this year, the page restrictions on references included in articles at the HCI field’s top conference, the ACM Conference on Human Factors in Computing Systems (CHI), were lifted. This policy decision had a profound impact on the citation practices of the CHI community. Our analysis of 11,542 articles published at the CHI Conference between 1981 and 2024 provides evidence of a year-by-year increase in the mean number of references included in CHI articles. This change signifies a departure from previous patterns and suggests the community’s entry into a novel trajectory (cf. Figure 1). If this trend continues undisrupted, an article in CHI 2030 will include on average about 130 references. Such a scenario poses practical challenges for authors and peer reviewers alike, straining the authors’ capacity to thoroughly engage with the literature and the peer reviewers’ capacity to vet the cited literature during the peer review process. It also raises fundamental questions about the nature of academic discourse and the balance between comprehensiveness and focus in scholarly communication.
The excessive and escalating number of references in CHI articles raises important questions about the sustainability of current citation practices. The challenges brought by the citation practices within the HCI community are multifaceted, encompassing the expansion of the conference itself and the elevation of standards for what constitutes a publication-worthy contribution. Such developments, while indicative of progress, also contribute to the increasing complexity of the academic ecosystem, negatively impacting both authors and peer reviewers. With the volume of references expanding, the feasibility of conducting thorough and effective peer reviews is called into question, suggesting a pivotal moment for reflection and reevaluation of how the HCI field approaches the foundational practice of citation. In light of these findings, this paper aims to spark a conversation within the CHI community and beyond about the future of scholarly communication and academic publishing. By examining the contextual factors and ramifications of changing citation practices, we invite scholars, practitioners, and policymakers to reflect on past decisions and identify sustainable pathways forward.
2. Related Work
2.1. The Field of Human-Computer Interaction
Human-Computer Interaction is both a field and a discipline (Liu et al., 2014; Blackwell, 2015). As a field, HCI encompasses diverse disciplines, such as computer science, cognitive science, psychology, design, sociology, anthropology, and more (Kim, 1995; Blackwell, 2015; Liu et al., 2014; Oppenlaender and Hämäläinen, 2023). HCI draws from these diverse areas to understand how humans interact with computers and how to design user-friendly systems. As an academic discipline, HCI is taught in structured curricula at many universities in dedicated degree programs, educating students in the principles and practices of HCI. Human-Computer Interaction is also recognized as a professional discipline with specific roles, such as UX researchers and interaction designers. Other researchers have aimed to look past the dichotomy of field versus discipline by framing HCI as problem-solving (Oulasvirta and Hornbæk, 2016). In the remainder of this paper, we refer to HCI as a field.
Each year, a growing number of articles are being published in the field of HCI. The international top conference in the field of HCI is the ACM CHI Conference on Human Factors in Computing Systems (CHI), which has been held each year since 1981 (with exception of 1984). CHI is the most prestigious conference venue for researchers in HCI, with many HCI labs submitting their articles only to this conference venue. As such, the CHI conference can serve as a reference standard for the entire field and discipline of HCI field. From its humble beginnings in the early 1980s, the conference has been expanding since 2005 in terms of articles accepted at the conference (see Figure 2). Except for the COVID-19 pandemic years 2021 and 2022, the CHI proceedings have exhibited near linear growth since 2005. In 2024, the CHI Proceedings exceeded 1000 articles. While this expansion is plan-driven and directed by CHI’s Steering Committee, it also reflects the expansion of the HCI field.
Figure 2. Number of research articles published in the CHI proceedings in the years 1981 to 2024 (excluding session details, panel sessions, and abstract-only entries).
[…]
Number of CHI articles from 1981 to 2024.
In this expansive and evolving research landscape, the practice of citing relevant works is important for HCI researchers. Citations serve as a vital link, connecting current research efforts with the rich history of past works. Citations acknowledge the contributions of other researchers and weave a rich tapestry of knowledge that informs and enriches new inquiries. For a field as interdisciplinary as HCI, where the integration of diverse perspectives and methodologies is crucial, citations provide a structured way to navigate and summarize the vast array of existing knowledge for readers without deep subject-matter knowledge and expertise. Further, citations enhance the credibility and rigor of academic discourse, enabling scholars to build upon a verified body of work. Proper citation practices contribute to the integrity of the HCI field, fostering an environment of collaboration and continuous learning. As such, citations are not merely a scholarly obligation but a critical element that underpins the academic and intellectual growth of the HCI community.
2.2. Meta-Research on the CHI Community and the Field of HCI
While the majority of HCI researchers conduct research within their field, some researchers have also examined aspects about the field of HCI. This is commonly referred to as meta-research (Ioannidis et al., 2015). Meta-research is valuable, as it enables the research community to reflect on its research practices (Pochat, 2024). The CHI proceedings, in particular, are a suitable and fruitful study subject for meta-scientific and bibliometric inquiries in the field of HCI. Bartneck and Hu, for instance, presented a scientometric analysis of the CHI proceedings, focusing on organizations and countries that contribute to CHI (Bartneck and Hu, 2009). Their work highlighted the difficulty of judging quality in the context of best paper awards, finding a mismatch between awarded papers and citations received. Our paper also includes an analysis of awards, however we focus on analyzing whether there is a systematic bias in awarded papers.
Lee et al. presented a bibliometric analysis of the CHI proceedings (1981 to 2018, excluding 1984), using a citation network analysis to identify emerging topics in CHI (Lee et al., 2019). Liu et al. also conducted a bibliometric analysis of CHI publications. Their keyword co-word analysis quantified and described the thematic evolution of the HCI field based on 3152 CHI articles published between 1994 and 2013. Like our work, their work includes a comparison between two time periods (1994–2003, 2004–2013), highlighting the underlying trends in the HCI community.
Pohl and Mottelson presented a quantitative meta-research study on paper writing at CHI, including an investigation on citations received. Their analysis of 6578 CHI papers found that readability, title length, and novelty influence citation counts. Our work differs in that we do not investigate writing, and only focus on citations given, not received. Another area of meta-research is the use of literature reviews in HCI. Stefanidi et al. conducted an analysis of CHI proceedings and ACM Transactions on Computer—Human Interaction (TOCHI) articles included 189 articles (Stefanidi et al., 2023). Like in our work, the authors note an “insufficient consensus of what to expect of literature reviews in HCI.” The diverse types and methodologies of literature reviews in HCI prompt us to split our analysis of literature reviews into a broad and narrow part.
Kaye presented a quantitative analysis of CHI, examining author counts, gender, and repeat authorship (Kaye, 2009). Among other findings, the work identified a trend toward an increasing number of authors per paper. We identify the same trend in our work, and similar to Kaye, we aim to encourage a discussion about about what the preferred state of CHI should be.
3. Method
3.1. Dataset Collection
The ACM CHI Conference on Human Factors in Computing Systems is the HCI field’s largest and most important annual conference. Therefore, the CHI conference proceedings are a fertile ground for investigating the citation practices of HCI authors. The CHI proceedings are stored in ACM’s Digital Library (ACM–DL), which is the main data source in our research.
We collected data on the number of references for each article published in the CHI proceedings from 1981 to 2024111With exception of the year 1984 in which the CHI conference was not scheduled. by scraping the ACM–DL. We excluded session details, panel sessions, keynotes, and abstract-only entries using regular expressions. For articles with more than 150 references, we manually collected the number of references due to the ACM–DL’s display limit. Additionally, we identified and corrected six instances where the number of references was incorrectly reported as zero. This process resulted in a dataset of DOI identifiers with their respective reference counts. We then downloaded the PDF documents for each DOI from the ACM–DL, resulting in a comprehensive dataset of 11,542 CHI articles.
4. Findings
Over the course of 34 years (from 1981 to 2024), authors at the CHI conference cited a total of 558,080 articles. The mean number of references per CHI article over the entirety of the CHI proceedings is 48.35 references ($SD=35.09$, $Min=0$, $Max=496$). However, the field of HCI has evolved from its humble beginnings, and a more in-depth look at these statistics is warranted. In particular, we focus on the rising number of mean references per article in the following section.
4.1. The Rising Number of References in CHI Articles
We first provide an historical overview of citation trends in the CHI proceedings, from the conference’s beginning to the current trend. We then extrapolate this trend into the future to provide an outlook for the year 2030. This year is not too distant (only six years from the time of writing), and we believe focusing on the year 2030 in our prediction of future trends will – if the current trend is not broken – provide an accurate and realistic estimation of the future of citations at CHI.
4.1.1. Historical development
In the early years of the CHI conference, it was not uncommon for authors to submit an article without a single citation to other works. The trend of zero-reference articles continued with only few exceptions until the year 1997 (see Figure 1 and Figure 3). From 1998 to 2017, the minimum number of references per article has been positive and in the one-digit range. Since 2017, the minimum number of references has increased into the double-digits.
Figure 3. Number of references per article in the CHI conference proceedings from 1981 to 2024.
[…]
Number of references per CHI article from 1981 to 2024.
Between 1981 and 2015, the mean number of references per CHI article showed only modest linear growth (cf. Figure 1). After 2015, a clear and sudden quantitative change in the CHI community’s citation practices took place (see Figure 1 and Figure 3). Clearly, the CHI community entered a novel trajectory at this point in time, characterized by a year-by-year increase in the mean number of references per article. While there were on average 28.8 references per article in the 2015 proceedings, the number of references per article more than tripled to an average of 88.6 references per article in 2024.
For about two decades since its inception, the maximum number of references per CHI article hovered at around 40. After 2015, the maximum number of references skyrocketed to up to 496 references in one article in 2022. The box plots in Figure 3 demonstrate that the maximum number of references per article has become more volatile in recent years. During the years of the COVID-19 pandemic (2020 and 2021), one can observe a slump in the maximum number and upper quartile of the number of references per CHI article. However, there is overall a clear tendency for CHI authors to include more references in their articles than before the year 2015, and this number is monotonously growing each year. The number of references per article between 2006 and 2015 can be approximated with linear regression to a function $y_{t+1}={m_{1}}y_{t}+b_{1}$, with $m_{1}=0.79$, $b_{1}=-1555.92$, and $y_{t}\in[2006,2015]$, with mean absolute error $MAE=0.50$ and mean squared error $MSE=0.38$. The trend after 2015 is approximated by the function $y_{t+1}={m_{2}}y_{t}+b_{2}$, with $m_{2}=6.68$, $c=-13435.27$, and $y_{t}>2015$ (with $MAE=1.88$ and $MSE=6.16$).
To examine the differences between the linear regression models fitted to the two distinct periods, we employed a Chow Test, a test to determine whether there is a significant difference in the coefficients between two the linear regressions. The Chow Test indicates a substantial divergence in the regression parameters across the two periods under consideration ($F=145.47$, $p=4.30×e^{-10}$). This extremely low p-value strongly refutes the null hypothesis, suggesting that the regression parameters (slope and intercept) for the two periods are significantly distinct. This statistical evidence underscores the presence of a notable change in the trend between the two periods, corroborated by a clear shift in the slope. We assert that the observed variations in the dataset characteristics are not merely coincidental but indicative of a fundamental shift in the underlying dynamics across the two periods examined.
4.1.2. Extrapolation of the observed trend
We extrapolate the observed trend with the linear regression model and predict the mean number of references per article in future instances of the CHI conference proceedings (see Figure 4). We estimate that if the current observed trend is not broken, an article published at CHI 2030 will contain on average 129.5 references. The minimum number of references per article in 2030 is predicted to reach 33.7 references, which is higher than the mean number of references from the year 2015 (28.9 references).
Figure 4. Extrapolation of the trend in the mean and minimum number of references per CHI article. CHI articles are predicted to include, on average, almost 130 references in the year 2030. The minimum number of references per CHI article is estimated to reach 33.7 references per article, which is more than the 2015 average level of references.
[…]
Prediction of mean and minimum number of references per CHI article from 2024 to 2030.
4.2. Root Cause: A Change in Editorial Policies at the CHI Conference
A diverse and complex community consisting of thousands of members, such as the CHI community, is not likely to change its course suddenly, unless change is mandated from top down. In this section, we discuss the editorial change that can be considered the root cause for the observed trend.
2015 CfP (excerpt)
A Paper is no more than 10 pages long, while a Note is no more than 4 pages long. This includes figures, references, appendices and an abstract of less than 150 words long. Over length submissions will be rejected.
2016 CfP (excerpt)
A Paper is no more than 10 pages long, while a Note is no more than 4 pages long. References do not count toward these lengths. The lengths do include figures, appendices, and an abstract of less than 150 words. Submissions that exceed these limits will be rejected.
Figure 5. Side-by-side comparison of excerpts from the CHI conference Call for Papers (CfPs) in 2015 and 2016 (own highlighting).
[…]
Side-by-side comparison of excerpts from the CHI conference Call for Papers (CfPs) in the years 2015 and 2016.
4.2.1. Lifting the page restrictions in 2016 launched the CHI community on its expansive path
The CHI conference organizing committee (Program Chairs) in consultation with the Executive Committee of the ACM SIGCHI (Special Interest Group on Computer-Human Interaction) introduced a decisive change in the year 2016. In this year, the conference’s Call for Papers (CfP) moved away from counting references towards the page count limit, as highlighted in the excerpts reproduced in Figure 5. Articles published at CHI ’16 were allowed to include an unlimited number of references, as long as the article’s content fit the 10-page limit, with references no longer counting towards this page limit. While sub-committees were introduced as another significant change in the year 2016, allowing an unlimited number of references clearly is one of the main contributors to the increase in the mean number of references identified in our work. Further, the 2016 submission guide for CHI authors stressed the importance of citing relevant previous work, explicitly asking “is prior work adequately reviewed?” The 2016 guide highlighted the importance of reproducibility of authors work, with “letting others build on your work” being regarded as “the entire purpose of a CHI Paper.” These direct quotes from the 2016 guide were stark reminders and may have contributed to making the community pay special attention to their references in the year 2016. These policy changes nudged the community towards its expansive path. Awards presented to articles may have also played a role, as demonstrated in the following section.
4.2.2. Are high-citation articles rewarded in the CHI community?
Each year, awards are given to a select number of accepted articles at the CHI conference. Typically, between 14.6% and 24.6% of articles are given an award at the CHI conference. [^2] Based on awards presented from 2015–2024. These awards are a signal for the community on which articles are to be considered high quality. As a quality signal, awards can motivate authors to change their behavior. If articles with a high number of references were to systematically be presented with awards, this could potentially incentivize CHI authors to include more references in their articles. In this section, we examine whether there is such a systematic bias in awarded papers. We analyze the difference in the mean number of references in articles that received an award at the CHI conference compared to the mean number of references in articles that did not receive an award. As award, we consider both “best paper” awards and “honorable mentions.”
Table 1. Comparison of mean number of references in articles awarded a best paper award or honorable mention (Mawarded) with the mean number of references in unawarded articles (Mregular) in the CHI Proceedings from 2015 to 2024.
Mawarded Mregular t p
2015 31.2185 28.1534 2.2984 0.0220 *
2016 43.4696 40.7047 1.6156 0.1068
2017 49.8843 44.4969 2.7794 0.0056 **
2018 57.8651 49.8795 3.6289 0.0003 ***
2019 61.1419 54.9910 2.6155 0.0091 **
2020 66.0513 60.1246 2.6780 0.0076 **
2021 72.9648 70.9868 0.6910 0.4898
2022 87.9219 81.9332 1.5201 0.1290
2023 88.9688 84.7736 1.2218 0.2221
2024 92.9579 87.6505 1.7950 0.0729
Table 1 shows no evidence of a systematic bias towards awarding high-reference articles in recent proceedings. For instance, in the 2024 CHI Proceedings, the mean number of references for awarded articles was 92.96 versus 87.65 for articles without award. This means a high number of references is not considered a signal of high quality in the CHI community in recent years. However, between 2015 and 2020 (with exception of 2016), awarded articles had significantly more references than unawarded articles. The systematic award of articles with above-average references may have sent a quality signal to authors, thus potentially contributing to launching the CHI community on its current trajectory.
4.3. Additional Contextual Factors influencing HCI Citation Practices
The CHI community is vast and highly diverse, and there are many more factors that potentially shape and make up the citation practices of the members of the CHI community. While the editorial change introduced in 2016 is the primary reason for the change in the citation practices of HCI scholars, in this section, we take a step back and shed light on a number of broader contextual factors affecting the citation practices at the CHI conference. We do not claim that this list of factors is exhaustive. Beside the investigated changes, other latent factors could contribute to the observed trend, as is the case with any complex system under change. The investigated factors are an increase in the average number of authors (Section 4.3.1), an increase in the number of literature reviews (Section 4.3.2), an increasing number of citations to arXiv pre-print articles (Section 4.3.3), and an increase in citations to data and code repositories (Section 4.3.4).
4.3.1. Collaborative research is increasing
There has been an increase in collaborative research efforts in the field of HCI, leading to articles with a larger number of co-authors. Figure 6 depicts the number of authors per CHI article for each CHI proceeding. The mean number of authors per article has increased over time, from an average of 1.85 authors per CHI article in 1981 to an average of 5.17 authors per article in 2024.
Figure 6. Violin plot illustrating the distribution and central tendencies (mean and median) of the number of authors per CHI article from 1981 to 2024, highlighting the trend toward an increase in collaboration over time.
[…]
Number of authors per CHI article from 1981 to 2024.
Figure 7. Mean number of references per author in CHI articles from 1981 to 2024.
[…]
Mean number of references per author in CHI articles from 1981 to 2024.
Figure 7 plots the mean number of references per author in CHI articles since the inception of the CHI Conference. In 1981, a CHI article would include on average 5.6 references per author. In 2024, this number has increased to 21.4 references per author. We note that while the mean number of references per author has seen a slow increase over time, there is a significant change in trajectory in the mean number of references per author after 2015.
4.3.2. Literature reviews are becoming more common
Literature reviews aim to capture the current state of a research area or field. Given that the production of scholarly articles is growing, we expect CHI authors are increasingly conducting literature reviews to address this expansion. We identified literature reviews by searching each CHI article’s full text (excluding the references section) for the keywords “literature review,” “literature survey,” and “systematic literature review.” We manually verified these occurrences, ensuring that only articles explicitly stating they conducted a literature review were included. Note that CHI authors, particularly in earlier proceedings, often refer to the related work section as a literature review. Additionally, some authors claim to have conducted an “extensive” or “exhaustive” literature review without detailing the process. Our data encompasses “literature review” in a broad sense, while “systematic literature reviews” are represented more narrowly, typically indicating articles that adhere to a more rigorous academic process.
Figure 8. Absolute number (top) and relative percentage (bottom) of articles in the CHI proceedings reporting having conducted a “literature review” (including “literature surveys” and “systematic literature reviews”; in red) and a “systematic literature review” (in teal color) from 1981 to 2024. Systematic literature reviews in CHI articles are a new phenomenon starting in 2016.
[…]
A line plot of literature reviews in CHI proceedings from 1981 to 2024.
In total, we identified 615 literature reviews and 60 systematic literature reviews. The time series in Figure 8 demonstrates that it is increasingly more common for CHI authors to conduct literature reviews. The amount of articles which mention having conducted a literature review was about 10% in the years 2021, 2022, and 2023. Figure 8 also demonstrates that systematic literature reviews in CHI are a recent phenomenon. Before 2016, there was no instance of authors calling their work a systematic literature review, but it has become increasingly more common for CHI authors to conduct systematic literature reviews. In 2024, the CHI proceedings included 17 systematic reviews (1.6% of the articles published in that year).
4.3.3. Pre-print articles are increasingly being cited by CHI authors
The landscape of academic publishing is changing. In this section, we examine how CHI authors cite pre-print articles. Pre-print articles are versions of scholarly articles that precede publication in peer-reviewed journals and conference venues. While traditionally shunned upon due to not being formally reviewed, it has become common place in some research fields, such as Machine Learning, to upload and also cite pre-print articles. Pre-print archives offer a convenient way for authors to claim priority capital, establishing the authors’ precedence in scientific discoveries or novel ideas (Newman, 2009). One of the most popular pre-print archives is arXiv, launched by Cornell University in 1991. In recent years, an exponentially growing number of articles are being submitted to this pre-print archive. The high number of available pre-prints and the success of pre-prints in certain fields, such as Machine Learning, may incentivize CHI authors to include more citations to pre-prints in their CHI articles. To investigate this in more depth, we scraped the references from the ACM–DL for each article in the CHI proceedings. We then identified the number of citations to arXiv by counting the number of times “arxiv” is mentioned in the articles’ reference section. For counting the number of occurrences of arXiv in the references section, we used a case-insensitive regular expression that accounts for hyphenation within the word arXiv due to line breaks.
[…]
Number of CHI articles citing arXiv pre-print since 2006.
Figure 9. Citations to pre-print arXiv papers (a) and percentage of CHI articles citing arXiv pre-prints $n$ times (b) in the CHI Proceedings.
Even though arXiv was made available to the World Wide Web in 1993, we find that none of the CHI articles before the year 2006 include citations to arXiv pre-print articles. 9(a) depicts the number of citations to arXiv pre-prints in CHI articles since 2006. We note there is a change in this time series after 2015. While the mean number of citations to arXiv pre-print articles has only slightly increased since 2016 (from 0.03 references in 2015 to 4.24 references in 2024), citations to pre-print articles have become overall more accepted among some CHI authors (9(a)). One article in CHI ’24 included 89 references to arXiv. 9(b) depicts the number of times $n$ that an article cites arXiv articles. From this figure, one can note an increase in the number of articles that cite a high amount of arXiv pre-prints. In recent proceedings, over a third of the CHI papers include more than one citation to arXiv. Given that this change coincides with the year 2016, citations to arXiv articles are one contributing factor to the increase in references per CHI article.
4.3.4. Data-driven research is on the rise
With the rise of big data and data-driven research methodologies, CHI papers may have started referencing more data sources, scientific software and tools, code repositories, and prior data-heavy studies. We investigate this trend by counting the number of references to five common open data and source code repositories:
  • osf.io, a cloud-based management solution for open access science, often used to share datasets and other data related to HCI studies,
  • zenodo.org, an open repository for datasets and other research related digital artefacts,
  • github.com, a source code repository that, with git LFS, can also store big datasets,
  • kaggle.com, a web-based platform for sharing datasets for machine learning, and
  • huggingface.com, a repository for storing and sharing machine learning related files.
As above, we limit our investigation to the articles’ reference sections which we scraped from the ACM–DL. [^3] We ignore citations in footnotes within the article’s full text to make the results comparable to the other findings in this article.
[…]
Mean number of citations to data and code repositories per CHI article.
[…]
Relative number of CHI articles citing data and code repositories at least once.
Figure 10. Mean number of citations to code and data repositories (a) and relative number of CHI articles citing at least one repository (b) in the CHI proceedings from 2008 to 2024.
[…]
Mean number of citations to data and code repositories (a) and relative number of CHI articles citing at least one repository (b) in the CHI proceedings from 2008 to 2024.
10(a) plots the mean number of references to code and data respitories in CHI articles. While the average number of citations to such repositories per CHI article is still low, we identify a clear increase in the number of citations to code and data repositories occurring after the year 2015. Before 2015, almost no authors formally cited these repositories, whereas today, such citations have become more common. In the proceedings of CHI 2024, 147 articles (13.9%) cite at least one data and code repository (see 10(b)). GitHub was cited most often with 115 articles (10.8%) in the CHI 2024 proceedings including at least one reference to this repository. This empirical trend in the citation practices reflects the broader trend towards applying data-driven methodologies and machine learning in HCI research.
5. Discussion
This article presented meta-research on the field of HCI, providing evidence of clear quantitative and qualitative changes in the citation practices of the CHI community co-occurring since the year 2016. Specifically, we have described five co-occurring changes:
  1. (1)
    An increase in the mean number of references included in CHI articles (Section 4.1):
    In the 2016 Call for Papers, references were no longer included in the page limit of CHI articles. This editorial decision launched the CHI community on an expansive path, characterized by a year-by-year expansion of the mean number of references included in CHI articles. One could argue that the observed trend in Figure 1 is simply an effect of the overall growth of publications in the HCI field. However, while growth may certainly be a contributing factor, the growth of the CHI conference proceedings precedes the observed trend by 10 years, as evident in Figure 2. Therefore, the increase in the mean number of references per CHI article is likely not a mere side-effect of CHI proceedings growth. Rather, it is the top-down policy change that took place in the year 2016 that opened the floodgates, allowing authors to cite more and more works in their articles.
  2. (2)
    An increase in collaboration and the number of references per author (Section 4.3.1):
    It is becoming more common for authors in HCI to collaborate and the mean number of authors per CHI article is expanding. An increase in the number of co-authors means that the set of authors may potentially contain authors from different disciplines. This implies that authors potentially draw from a broader range of interdisciplinary sources, hence leading to more citations to other articles. However, our analysis found that the increase in the mean number of authors is not proportionate to the increase in the mean number of references, as evident in Figure 7. The growth in the mean number of references exceeds the growth in authors.
  3. (3)
    An increase in the number of literature reviews (Section 4.3.2):
    The literature is growing at a rapid rate, and we found that CHI authors increasingly conduct systematic literature reviews of the expanding corpus of scholarly literature. The COVID-19 pandemic years saw an acceleration of this trend, denoted by an increase in the relative number of CHI articles conducting literature reviews. This can be explained by authors not being able to conduct user studies, a common method in HCI, during the pandemic. However, as depicted at the top in Figure 8, the growing trend toward literature reviews already started earlier. The editorial decision to lift the page restrictions greatly contributed to the growing popularity of literature reviews at CHI. Literature reviews typically contain an above average number of references, which contributes to the increase of the mean number of references included in CHI articles.
  4. (4)
    An increase in the number of citations to arXiv (Section 4.3.3):
    Entire fields of science rely on arXiv as a means for article dissemination. The field of HCI is not immune to this trend, and it is becoming increasingly more common for CHI authors to cite unrefereed articles. Before 2015, citations to arXiv were shunned upon by CHI authors. From 2016 onward, citations to arXiv were no longer competing for space with citations to peer-reviewed articles. The editorial decision of lifting page restrictions in CHI articles contributes to this shift in citation behavior. Today, many CHI authors cite arXiv pre-print articles. This can lead to adverse problems, such as the pre-print being cited instead of the published article.
  5. (5)
    An increase in the number of citations to open data and code repositories (Section 4.3.4):
    Before 2015, there were virtually no formal citations to code or data repositories to be found in the references sections of CHI articles. From our analysis, it can be concluded that lifting the page restrictions animated authors to include more citations to such repositories in their CHI articles. From an economic lens, there no longer is a marginal cost to including another citation to arXiv or code and data repositories and there is no trade-off cost between citations.
Together, these co-occurring trends shape the citation practices of the CHI community in the shifting landscape of HCI. These trends reflect an interesting change of pace in academic working in the CHI community starting with CHI ’16. The trends are also a reflection of broader changes that affect not only the field of HCI, but academia in general.
What is striking about the observed trend in Figure 1 is that it follows a relatively stable linear growth pattern. This near-linear growth takes place in the presence of exponential growth in academia, such as the exponentially growing number of scholarly articles. Future work could investigate why the observed year-by-year growth in references is linear. Perhaps it could be explained with a simple psychology model of herd mentality. Each year, members of the CHI community observe — whether consciously or sub-consciously — an increase in the mean number of references per CHI article. In other fields of science, evidence has been found that longer articles are cited more (Kousha and Thelwall, 2022). A simple strategy of “the more, the better” could be motivating authors in CHI to include more references in their articles to satisfy the seemingly growing expectations of the peer reviewers. In the following section, we discuss why the growing number of references is a critical concern that demands attention.
5.1. Implications for Authors, Reviewers, and the Broader Academic Community
An article with many references can help novices and those unfamiliar with a research area to learn and become familiar with the research area. However, we argue the observed trend (i.e., excessive citations as evident in the rising number of references in CHI articles) is a negative development for authors, peer reviewers, and the entire field of HCI, for a number of different reasons.
First, we argue that 130 references per article is too many to thoroughly screen during the peer review process. A high number of references increases the cognitive load on peer reviewers, contributing to reviewer fatigue. This only adds to the already high amount of peer review fatigue, with conferences struggling to recruit high-quality reviewers. Peer review is the backbone of academic rigor, and if reviewers only give a cursory glance at the reference section, the integrity of the peer review process is compromised.
Second, the number of references included in CHI articles serves as a signal to the CHI community. Peer pressure and herd mentality suggest that if the average number of references is high, authors feel compelled to include a similar number of references in their articles. Deviations from this norm are likely to be viewed negatively by peer reviewers, creating a spiraling pattern of year-by-year increasing academic pressure and competition. This spiraling pattern contributes to author dissatisfaction and fatigue.
Third, despite the shift away from awarding best paper awards and honorable mentions to high-reference articles, the overall rise in the number of references still sends the wrong message to junior members of the CHI community. Today, PhD students in HCI are faced with a seemingly insurmountable and overwhelming amount of literature work, which distracts them from focusing on developing their own research skills and expertise.
Fourth, the inflation of citations diminishes the value of each citation. A citation today does not carry the same weight it did in the past. This is similar to how early search engine optimization efforts devalued the significance of web links. The increasing number of citations reduces their impact and can lead to an overload of less meaningful references.
Last, the high number of references makes it easier for bad actors to commit fraud. In an environment of hypercompetition and “perverse” incentives (Edwards and Roy, 2017), academic fraud is surprisingly common (Tijdink et al., 2014). With no clear trade-off cost between citations in an article, the reference section can be expanded arbitrarily, facilitating citation manipulation. This can include self-citation (Bartneck and Kokkelmans, 2011), author collusion in citation cartels (Kojaku et al., 2021; Jolly et al., 2020), and even the purchasing of citations (Ibrahim et al., 2024). The high number of references that, likely, undergo only cursory peer review, makes it harder to detect such unethical and fraudulent research practices.
5.2. Reflective Considerations for Future Research Practices
Many of the developments in the CHI community are plan-driven. For instance, the number of articles accepted at the CHI conference is deliberately expanded each year. However, one should not forget that top-down policy changes lead to behavioral changes in members of the research community. Notably, while the CHI Steering Committee initially started the CHI community on its expansive path, it is the collective decision of authors at CHI to, year-by-year, cite more and more articles. In this section, we urge the community to pause and reflect on its citation practices.
This work presented trends in citation practices to foster dialogue — not to prescribe mandates, but to encourage diverse interpretations and applications of these findings. We invite readers to engage with the data, reflecting on its implications for the future of HCI research. Our data serves not as a directive, but as a mirror reflecting the current state of CHI’s scholarly practices. By offering this comprehensive overview, we aim to help scholars and practitioners navigate the complexities of the modern research landscape.
However, a critical question arises: How much responsibility does each author in the CHI community bear for the observed changes in citation practices? While scientific knowledge advances rapidly, the academic system remains relatively conservative and resistant to change. This system, dominated by powerful publishers like ACM Press, often feels too vast for any one individual to influence significantly. Therefore, we call for reflective consideration. We urge CHI authors to be more deliberate with their citations, viewing them as quality signals and endorsements. Readers should consider how these trends impact their research practices and the broader HCI community. Through introspection, the community can shape future research in ways that are thoughtful and aligned with shared values. This approach emphasizes the agency of the scholarly community to enact change. It respects the diversity of perspectives within our field and acknowledges each researcher’s autonomy to interpret and integrate these insights according to their goals.
5.3. Potential Solutions
One potential response to the observed trend is to simply ignore it, though this approach risks future challenges. Proactively addressing the issue of an increasing number of references is a more constructive approach. This section discusses two viable solutions for addressing the growing number of references.
5.3.1. Solution 1
A first solution would be to revert to the 2015 editorial policy (or an updated version thereof). This could involve setting a limit on the number of references or the overall number of pages per article, as was the case in 2015. At that time, citing a paper was an endorsement, and any citation had to be weighed against the space it would take away from the article’s content or other references. This trade-off has since been lost. While modern practice tends to allow an unrestricted number of pages and references, recent iterations of the CHI Conference have seen inconsistently enforced page limits. For instance, the quality signaling by the Chairs during the CHI ’23 submission process clearly indicated that unrestricted growth in article length is not welcomed by the Conference Chairs and the Steering Committee. Limiting the number of references for most articles (with the exception of literature reviews) would be a viable solution to address the growing and excessive number of references included in CHI articles.
5.3.2. Solution 2
A second potential solution pertains to the nature of the peer review process in academia. Peer review is the cornerstone of scholarly rigor, serving as a crucial filter that ensures the quality, relevance, and integrity of academic publications. From the perspective of creativity theory (Runco and Jaeger, 2012) and Csikszentmihalyi’s systems model of creativity (Csikszentmihalyi, 2014), the peer review process is a gatekeeping process that validates two critical components of an article: Novelty and usefulness (or appropriateness). As gatekeepers, it is the peer reviewer’s task to assess these two components and, hence, the creativity of a submitted article. In some cases, an extensive related work section may support the evaluation of an article’s novelty, especially in cases where the peer reviewer may not have sufficient knowledge for conducting the review. In an ideal world, a peer reviewer lacking sufficient knowledge of the field would not review an article in the first place. However, in practice, a mismatch between reviewer expertise and with a paper is, unfortunately, rather common (Shah, 2022), especially in large and highly diverse conferences, such as CHI. However, in the role of a gatekeeper, it is the peer reviewer’s duty to be knowledgeable about works in the field. However, in a field as vast and diverse as HCI, this can be challenging.
A potential solution is to split peer review into two parts, one conducted by the human peer reviewer, the other one performed by machine learning. Language models, as knowledge compressors, can assess whether a contribution is original and whether the references are comprehensive more effectively than human reviewers. This would allow human reviewers to focus on assessing the second component — the usefulness and appropriateness of the methodologies, findings, and contributions.
Evidence suggests we are moving towards this direction. In the United States, automated essay scoring systems are already grading student essays (Feathers, 2019; Weatherbed, 2024), and a growing number of peer reviewers use language models to write their reviews. A 2024 study by Liang et al. found that up to 17% of peer reviews at top machine learning conferences are substantially authored by language models (Liang et al., 2024). Liang et al.’s study found an overlap of between 30.9% and 39.2% between the points raised by GPT-4 and human reviewers, and their user study with 308 researchers from 110 institutions in the US highlighted that researchers find AI-generated feedback useful (Liang et al., 2023). Similarly, Sun et al. found automated peer review significantly expedited the review process and improved review coverage, although raising concerns around trust, bias, and agency (Sun et al., 2024).
Automating aspects of peer review could offer significant benefits. The cost of peer review is estimated to be billions of dollars annually (Aczel et al., 2021). In a system strained by an overwhelming volume of literature (Hanson et al., 2023), partial automation can alleviate the burden on reviewers. Further, automating peer review could increase trust in the scientific process by reducing biases, such as the Matthew effect, which has been a concern in academia since 1968 (Merton, 1968). In the future, language models could assist in evaluating novelty, leaving the assessment of usefulness to human reviewers, thereby enhancing the efficiency and reliability of the peer review process.
6. Conclusion
This article presented meta-research on citation practices at the ACM CHI Conference, the top conference in the field of Human-Computer Interaction. Our segmented regression analysis provided clear evidence of a change of pace in academic working in the CHI community, indicative of broader trends in HCI. We provided empirical evidence of a year-by-year increase in number of references included in CHI articles. The field’s excessive citation practices were enabled by a policy decision that lifted page restrictions on the reference section of CHI articles. The implications of such meso-level policy decisions can be profound. This study highlights the value of meta-research for HCI and the CHI community. More meta-research is needed to carefully consider the broader impact of policy decisions and to provide clear recommendations for stakeholders. It is time for the HCI community to pause and reflect on its citation practices, and constructively discuss solutions. A different pathway forward is needed, in particular with an eye on the research community’s junior members and the sustainability of the HCI field.